Circular No. 8722

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

2002 JF_{56}

L. Barrera, Universidad Metropolitana de Ciencias y Educacion, Santiago; and C. Tubiana, R. Duffard, and H. Boehnhardt, Max Planck Institute for Solar System Research, Katlenburg-Lindau, write that minor planet 2002 JF_{56} was observed with the FORS2 instrument at the European Southern Observatory’s 8.2-m Very Large Telescope telescope ‘Antu’ on four nights between May 25 and June 2. Preliminary analysis of the visible spectra indicate an S-type taxonomy for the object, which is being viewed during a distant fly-by of the ‘New Horizons’ spacecraft on June 13.

COMETS C/2006 J9, C/2006 J10, C/2006 J11 (SOHO) Three additional Kreutz sungrazing comets have been found on SOHO

website images (cf. IAUC 8720). C/2006 J9, which was also found by T. Chen, was stellar and reached mag 7.3 on May 10.579 UT at 5.9R_{⊙}. C/2006 J10 was small and diffuse, reaching mag 7.8 on May 11.204 at 6.0R_{⊙}. C/2006 J11 was small, diffuse, and too faint for photometry.

Comet 2006 UT α_{2000} δ_{2000} Inst. F MPEC
C/2006 J9 May 10.564 3^{h}13.^{m}6 +16^{o}20′ C2 HS 2006-L32
C/2006 J10  11.204 3 15.3 +16 22 C2 TH 2006-L32
C/2006 J11  11.746 3 18.1 +16 34 C2 JS 2006-L32

COMET 176P/LINEAR = (118401) LINEAR The Committee on Small Bodies Nomenclature has agreed to name

the comet 1999 RE_{70} (cf. IAUC 8704), which has been assigned also the permanent comet number 176P (MPC 56919), with the same name as the minor planet (118401): LINEAR, for the ‘Lincoln Laboratory Near-Earth Asteroid Research’ survey that discovered the object (cf. MPC 56963).

SUPERNOVA 2006da Further to IAUC 8622, R. Quimby reports the discovery of an apparent

supernova in unfiltered CCD images taken on June 12.41 (at mag ∼ 16.6), and 13.41 UT (mag ∼ 16.8) using the 0.45-m ROTSE-IIIb telescope. SN 2006da is located at α = 23^{h}27^{m}48^{s}.57, δ = +14^{o}28′27′′.0 (equinox 2000.0), which is 1′′.5 west and 4′′.2 south of the center of the apparent host galaxy. Nothing was visible at this location in a co-addition of images taken between 2005 July 15 and 2005 Sept. 12 (limiting mag ∼ 18.8).

2006 June 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green